Status Report

NASA Space Station On-Orbit Status 16 Feb 2004

By SpaceRef Editor
February 16, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 16 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.  Underway: Week 17 of Increment 8.

After an early wakeup at 9pm EST last night (TsUP/Moscow time 5am this morning) and station inspection CDR/SO Michael Foale and FE Alexander Kaleri performed their 8th

session of the periodic Russian biomedical assessments Body Mass Measurement (PZEh-MO-8) and Calf Volume Measurement (PZEh-MO-7) before breakfast and first exercise.  Kaleri set up the MO-8 “scales” equipment and later broke it down and stowed it away. (Last time done: 2/3)   [For operational monitoring and data recording the crew used the special IFEP (in-flight examination program) software on the medical equipment computer (MEC).]

Soyuz thruster testing, Mike Foale performed the regular routine status checkup of the autonomous Increment 8 payloads in the Lab. 

Alex Kaleri later conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO).   [Maintenance generally consists of inspection of fluid lines/connections and filter replacement.]

Sasha also attended to his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.

The crew jointly reviewed the uplinked preliminary timeline for the EVA-9 on 2/26; they also studied the applicable Russian operations data files (RODFs), supported by tagup with ground specialists via S-band.   [During the spacewalk, the CDR will be EV1, the FE EV2.  Opening of the EV hatch and egress from the DC-1 will be at ~ 4:20pm, right after the first of four eclipses (i.e., orbital night periods, during which Orlan EVAs, as opposed to EMU EVAs, come to a stop for crew rest).  Estimated duration of the EVA is 5 hours 31 minutes.  The very busy timeline lists 46 individual activity steps, including translations to worksites.]

In further preparation for the Orlan spacewalk and the 2/19 training session, Kaleri and Foale gathered and laid out the replaceable components and auxiliary gear for the Orlan “skaphandr” suits, complete with fit checks.   [Assembled were portable O2 tanks (BK-3), storage batteries, LiOH canisters, moisture collectors, KVO liquid cooling garments, ShL-10 headsets, GP-10K gloves, BK-10 undergarments, socks, filters for FOR feedwater line, IK Orlan measurement unit and BOS degassing pump, etc.]

Thruster testing of the 7S/Soyuz TMA-3 MCS (motion control system, Russian: SUD) took place this morning after attitude control handover to the RS (Russian segment) at 3:30am EST.  The station was in free drift from 3:44am for about 15 minutes.  Return to US attitude control was at 4:10am.  Ten minutes later attitude was back in LVLH YVV mode (local vertical/local horizontal, y-axis in velocity vector), to be maintained until this weekend (2/21).

Mike worked briefly on the MSG (Microgravity Science Glovebox), tending the PromISS-3 (Protein Crystal Growth Monitoring by Digital Holographic Microscope 3) experiment by removing videotape #8 and installing tape #9.

Starting after lunch (5:15am), the crew dismantled the Kurs-A rendezvous and approach radar system of the 13P/Progress-260’s motion control and navigation system (SUDN), and removed it from the cargo drone, a three-hour job.  These valuable components, which include the BTsVK onboard digital computer, will be returned to Earth for “recycling” (when Shuttle flights have resumed).   [KURS-A is the active half of the Russian space program’s proven S-band radar system for automated flight, which measures relative motion parameters between Progress (or Soyuz) and the ISS during rendezvous operations, to enable the autopilot’s calculation of corrective impulses.  There are two sets of KURS-A electronics (for redundancy) housed in a common container in the Cargo Module.  The passive counterpart (KURS-P) of the system is on the Service Module (SM), with one antenna each at the tip of the two solar array wings.]

The Elektron oxygen generator continues to be off.  No SFOG (solid-fuel oxygen generator) “candle” was used as yet, and there is still O2 stored in 13P.

Yesterday, the crew observed another small (2-inch) piece of debris floating away from the ISS, apparently of Russian origin, from the 7S/Soyuz solar arrays.  [A possible second, washer-like object was noticed on one of the downlinked images.  A video was also recorded of the FOD (foreign object debris) using an SSRMS/robotarm camera.  Preliminary info from Moscow indicates that the eyebolt may be from the Soyuz solar arrays (S/A) where four of them are used to safe the S/A during launch with a hook mechanism which is released via pyro after insertion.  They are no longer needed after orbital insertion.  The bolts are secured with a nut and a locking wire, and apparently one of them came free.  Similar bolts are used on the Progress.  More at tomorrow’s IMMT (ISS Mission Management Team) meeting.]

The CDR reported being awakened by an “electric motor noise” in the vicinity of the S1-S2 racks at around 9:15pm EST last night.  At about that time, MCC-H commanded “Zoom Out” on the Lab starboard ETVCG (external TV camera group) camera, but its noise seemed to sound differently to Mike.  Under investigation.
FE Kaleri reported yesterday that during his evening exercise on the Russian VELO ergometer he noted that the frame underneath the generator became hot, to the point of being “too hot to touch”. He estimated the temp to be 70C.   Currently under review by TsUP/Moscow, but VELO ops will continue meanwhile.

Foale and Kaleri worked out according to their regular daily physical exercise program on TVIS treadmill and CEVIS cycle (aerobic), VELO ergometer with force loader and RED exerciser (anaerobic).

The crew’s sleep time began at 12:30pm EST, to extend until tonight, 9:00pm.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Nihoa island, Hawaiian chain (this mile-long island is the eroded summit of a volcano.  Images of the coastline are being requested), Shanghai, China (looking slightly left on the south side of the estuary), Internal waves, S China Sea (looking right towards the glint disc to document internal waves set up in the shallow seas around the bulge of Vietnam, due south to the island of Borneo.  Areas of internal waves are being mapped globally), Kabul, Afghanistan (nadir pass), Delhi, India (looking slightly left on the Yamuna River), Plankton, Oman (Dynamic event.  Blooms appear on SeaWiFS imagery in the Persian Gulf region.  On this descending pass, the crew was asked to look in the Persian Gulf [right of track], then in the Gulf of Oman [left of track] into the Arabian Sea), and Subglacial river courses, Libya (slightly sinuous patterns on dark rock surfaces are the courses of large ancient rivers that flowed under an ice cap.  A mapping swath [with a shorter lens for general orientation] parallel with track and left of track, was requested.  Visual cues for the site are (1) the dark color of the ancient rocks, and (2) the more distant Marzuq sand sea left of track.

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 3:00am EST, 2/13).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.5; ppO2 (mmHg) — 149.3; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.9.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 23.0.
  • Node:  Pressure (mmHg) — 738.16; temperature (deg C) — 22.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 740.57; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 740.0; temperature (deg C) — 23.2; shell heater temp (deg C) — 23.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.4
  • PMA-2:  Shell heater temp (deg C) — 10.8

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); battery #4 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 11:31am EST [= epoch]):

  • Mean altitude — 366.9 km
  • Apogee — 371.7km
  • Perigee — 362.1 km
  • Period — 91.88 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.0007073
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 110 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29925

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.